OpenStudy (anonymous):

Explain how to simplify x(2x + 3x2)

Parth (parthkohli):

You'd simplify it by distributive property.

OpenStudy (anonymous):

Distributive law

OpenStudy (anonymous):

a(b+c) = ab + ac

Parth (parthkohli):

\[x(2x + 3x^2) \implies x(2x) + x(3x^2) \]

Parth (parthkohli):

Or you could say\[a(b \pm c) = ab \pm ac \]

OpenStudy (anonymous):

x(2x + 3x2) = 2x^2 + 3x^3

OpenStudy (lgbasallote):

....or you take out another x

OpenStudy (mayankdevnani):



OpenStudy (mayankdevnani):

@sammiecupcake666420